MHATT-CAT Useful Operational Data for Run 3 of FY03
Eric Dufresne, Dohn Arms, Don Walko
MHATT-CAT
Started May. 13, 2003
(www.mhatt.aps.anl.gov/Sectors/Sector7/Operations/FY03_run3/)
Table of Content:
Introduction.
The work shown below is an account of useful operational data taken during
run 3 of FY03 on 7ID. This page will contain useful stability information on the
7ID High Heat Load Monochromator and also on the various repairs and problems
identified during the run.
Summary of the May 2003 shutdown activities.
- The APS User meeting was quite a success. See some pictures of a few
MHATT-CAT Users here
- The APS repaired several faulty O-ring on the Experimental Hall on the
hot water system. This affected all sectors. See this picture for a view near 7ID-D.
- The 7ID-C Kappa Y1 slider was repaired. See the full story here.
- Dohn Arms installed gas bottle mounts for the 7ID-B and C hutch.See some pictures here.
- Dohn Arms successfully upgraded the 7ID-D VME crate to the latest stable
release of EPICS. The other beamline crates will follow this year.
Several beamline computers were also protected by firewalls due to a new
ANL cybersecurity policy.
First day of beam time, May 28, 2003.
The run started on 5/28 without any significant delay. User beam was provided
nearly immediately this run because no commissioning time was used during this
start up. The 7ID-B mini-hutch needed a radiation survey since it was reopened
during the shutdown for the PSS validation.
Fig 05-31.1 show the intensity during a time series
started on 05/31/03 at 00h11 and lasting 62 hours or so. The beam is in non
top-up mode with the 324 bunch mode so the intensity varies following the ring
current. The ID gap was set to 12.05 keV, the mono to 12 keV, the white beam
slits to 0.35 mm by 0.35 mm.
Fig 05-31.2 shows the beam position measured with the
7ID-C X-ray BPM. The beam moves due to the LN2 fill and the ring fill-on-fill.
The typical beam motion due to LN2 fills are about every 4 hours with a 40 um
peak to peak variation.
Fig 05-31.1. Time series of the beam intensity in 7ID-C, starting at
00h11 on 05/31 and
lasting 62 hours. The 7ID-C diode sum, the ring current, ion chamber in B and C
are shown. The 7ID-C ion chamber is not connected.
Fig 05-31.2. The beam position, 49.2 m from the source, or 19 m from the
High Heat Load mono in 7ID-C during the same time series as Fig. 05-31.1.
Wedn. June 11, 2003.
Fig 06-11.1 shows the derivative spectrum of the Ta
absorption factor ln(I/I0) versus the monochromator energy. The monochromator
was off by a 2.5 eV (low side) and was recalibrated. The edge which is defined
as the peak of the derivative spectrum is at 9.881 keV. 0.5 eV steps were used
for the measurement.
Fig 06-11.1. The derivative spectrum of the Ta edge. Before and after the
calibration on 6/11/03.
Thursday June 12, 2003.
Fig 06-12.1 show the intensity during a time series
started on 06/12/03 at 01h43 and lasting 8 hours or so. The beam is in top-up
mode now. The ID gap was set to 10.055 keV, the mono to 10 keV, the white beam
slits to 0.35 mm (V) by 0.1 mm (H).
Fig 06-12.2 shows the beam position measured with the
7ID-C X-ray BPM. The beam moves due to the LN2 fill like it has since the
beginning of the year. The typical beam motion due to LN2 fills are about every
4 hours with a 40 um peak to peak variation.
Fig 06-12.1. Time series of the beam intensity in 7ID-C, starting at
01h43 on 06/12 and lasting 8 hours.
The 7ID-C diode sum, the ring current, ion chamber in B and C
are shown.
Fig 06-12.2. The beam position, 49.2 m from the source, or 19 m from the
High Heat Load mono in 7ID-C during the same time series as Fig. 06-12.1.
Thursday July 24, 2003.
Fig 07-24.1 show the intensity during a time series
started on 07/24/03 at 00h54 and lasting 96 hours or so. The beam is in top-up
mode now. The ID gap was set to 10.055 keV, the mono to 10 keV, the white beam
slits to 0.2 mm (V) by 0.2 mm (H).
Fig 07-24.2 shows the beam position measured with the
7ID-C X-ray BPM. The beam moves due to the LN2 fill like it has since the
beginning of the year.
Fig 07-24.1. Time series of the beam intensity in 7ID-C, starting at
00h54 on 07/24 and lasting 96 hours. The 7ID-C diode sum, the ring
current, ion chamber in B are shown.
Fig 07-24.2. The beam position, 49.2 m from the source, or 19 m from the
High Heat Load mono in 7ID-C during the same time series as Fig. 07-24.1.
Lessons learned during run 2 of FY03
Vacuum trips MF1+MF2 on 2/01/03
Vacuum trip MF2 on 2/08/03 10 am
Set point raised on MF2 on 2/12/03 to 2E-4 from 2E-5 Torr by ED.
Mono Huber tripped EPS on 3/12/03 when Don changed the mono energy.
Yoruba was suspended on 3/21/03 when someone logged out...
2nd time it happens
Dohn Arms fixed this for good.
Crocooler fill level: keep buffer between 50 and 60 %.
Vacuum trip MF1+MF2: on 4/10/03 1h50 pm. MF1 controller ultimately tripped.
Real gas load noticed on mono pumps when GV3 open.
Should we replace the pump?
To do list:
-check and fix Y-BPM motor/driver (see 1/29/03).
-fix the EPICS LN2 level sensors on the Oxford cryocooler.
-complete remote shutter interface
-Reduce the speed of the pressure response of the Huber chamber gauge.
-Remove the suspend option for the OS on yoruba and all SUNs... (Done)
-Prevent the turbo from venting when the pump is turned off or a power failure
occurs.
|